OpenStudy (anonymous):

Four friends were born on consecutive years. The sum of their ages is 62. What is the age of the youngest friend?

OpenStudy (shadowlegendx):

x + x + 1 + x + 2 + x + 3 = 62

OpenStudy (shadowlegendx):

4x + 6 = 62 Subtract 6 from both sides 4x = 56 Divide 4 from both side x = 14 14 + 14 + 1 + 14 + 2 + 14 + 3 = 62 62 = 62 The youngest age is 14

OpenStudy (anonymous):

hmm okay i understand i didn't even think of it like that

OpenStudy (shadowlegendx):

It's really easy c;